OpenStudy (anonymous):

solve for x log[3](x+3)-log[3](x-13)=2

OpenStudy (anonymous):

first write as \[\log_3(\frac{(x+3)}{(x-13)})=2\]

OpenStudy (anonymous):

then switch to exponential from via \[\frac{x+3}{x-13}=3^2=9\]

OpenStudy (anonymous):

so you get \[x+3=9(x-13)\] and you can take it from there

OpenStudy (anonymous):

x=15?

OpenStudy (anonymous):

yes 15 is it!